OpenStudy (anonymous):

Perform the indicated operation. Simplify if possible. \frac{ x^2-6x }{x-2}+\frac{ 8 }{ x-2 }

OpenStudy (anonymous):

\[\frac{ x ^{2}-6x }{ x-2 }+\frac{ 8 }{ x-2 }=\frac{ x ^{2}-6x+8 }{ x-2 }\] -6=-4-2 -4*-2=8 1*8=8 factorise the numerators and denominators and cancel the common factor.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\frac{ x ^{2}-4x-2x+8 }{ x-2}=\frac{ x \left( x-4 \right)-2\left( x-4 \right) }{ x-2 }\] \[=\frac{ \left( x-2 \right)\left( x-4 \right) }{x-2 }\] =x-4

OpenStudy (anonymous):

thanks I was confused